x^2 + 10x = 300
Find the positive value for x.
Give your answer correct to 3 significant figures.

Answer by Jk22(389)

x^2+10x-300=0
x=(-10+sqrt(100+1200))/2=-5+5*sqrt(13)=5(-1+sqrt(13)) (+ to get a positive ans)
Square root of 13 can be extracted manually :
13...|3,605
9----|
------------
40.0.|66x6=396
396..|
------------
40.0.|680x0
40000|6805x5

hence : x=5*2,605~13,025